There was a time when a man was twice the age of his wife but the next year he was only 1.5 years older. If he is 44 how old is his wife?
Answer by macston(5194)

.
x=husband's age when he was twice wife's age; y=wife.s age when husband was twice her age;
H=husband's current age; W=wife's current age
.
x=2y
x+1=1.5(y+1)
2y+1=1.5y+1.5
0.5y=0.5
y=1 Wife was one year old when husband was twice her age.
.
x=2y
x=2(1)=2 Husband was 2 years old when he was twice wife's age.
.
x-y=2-1=1 year Husband is one year older than wife.
.
H-1 year=W
44 years - 1 year=W
43 years=W
ANSWER: Wife is 43 years old.
.